# Suicide Risk Assessment on Social Media with Semi-Supervised Learning

### v1

@@ Abstract
Index Terms:
Natural Language Processing, Suicide Risk, Social Media, Semi-Supervised Learning
I Introduction
@@
II Materials and methods
II-A Task and Data Description
The Suicide Detection on Social Media BigData Cup Challenge  [^2] https://www3.cs.stonybrook.edu/~ieeebigdata2024/ involves creating a predictive model to classify Reddit posts into four suicide risk levels: 1) Indicator: The post content has no explicit expression concerning suicide; 2) Ideation: The post content has explicit suicidal expression but there is no plan to commit suicide; 3) Behavior: The post content has explicit suicidal expression and a plan to commit suicide or self-harming behaviors; and 4) Attempt: The post content has explicit expressions concerning historic suicide attempts.
As shown in Table I, the challenge organizers developed a corpus composed of 500 labeled and 1,500 unlabeled instances. A test dataset that includes 100 unlabeled instances is also included [12]. Each instance contains the text content of a Reddit post taken from the r/SuicideWatch sub-Reddit.
However, two major challenges are presented with this dataset. The first challenge is that only a small portion of available data is labeled, limiting the efficacy of a traditional supervised approach. The second challenge is that the intrinsic class-imbalance present in the labeled dataset makes it difficult for models trained on the provided data to predict classes underrepresented in the labeled corpus (i.e. Attempt).
TABLE I: Statistics of the Suicide Detection on Social Media dataset.
@@ II-B Method Development
II-B1 Backbone Model Selection
@@
II-B2 Semi-Supervised Learning with Stratified Confidence Sampling
@@ II-B3 Label Correction
II-C LLMs with Zero-Shot Learning and Fine-Tuning
@@
We evaluated the capability of Llama3-8B to detect suicide risk levels in the zero-shot learning setting and fine-tuning setting. For fine-tuning, we compared two strategies: next-token prediction and sequence classification. For both zero-shot learning and fine-tuning, the prompt starts with the definitions of the four suicide risk levels, followed by the input text (see Supplementary Table I).
II-D Evaluation Metrics
@@ II-E Experimental Settings
III Results
III-A RoBERTa Outperforms Other Models
Before evaluating our semi-supervised approach, we first fine-tuned three transformer-based models using a supervised approach to determine which embedding model performs the best and establish a baseline against which we can compare our semi-supervised methods. Three transformer-based models, BERT [16], BiomedBERT [17], and RoBERTa [14] were trained and validated according to our experimental settings outlined in Section II-E. Figure 2 shows the five-fold cross-validated performance of our systems. In columns labeled ”Indicator”, ”Ideation”, ”Behavior”, and ”Attempt”, we plot the binary-F1 score corresponding to posts of that class. The F1 scores yielded by our supervised models indicate that RoBERTa outperformed the other two embedding models across all F1 metrics (both class-specific and multi-class). This improvement is especially pronounced when examining the class-specific metrics. A more detailed report comparing the performance between the three models can be found in Supplementary Table II.
@@ III-A RoBERTa Outperforms Other Models
After deciding on RoBERTa as an embedding model, we conducted self-training with SCS pseudo-label acquisition to generate a dataset of 1,500 pseudo-labeled posts according to the methods outlined in Section II-B2. Furthermore, we also partially verified these pseudo-labels according to the strategy outlined in Section II-B3.
III-B Correcting Initial Pseudo-Labels
@@
After training five models and generating five sets of pseudo-labels, we calculated the agreement between each set of pseudo-labels and the majority-vote pseudo-labels using Cohen’s Kappa score [28] (Supplementary Table III). The rate of agreement between the five sets of pseudo-labels is high, which makes sense given that the training data between any two of the five models is 75% similar. To determine which samples to annotate, the final set of 1,500 pseudo-labeled posts was first split into four separate subsets based on the number of votes the majority voted class received between the 5 initial pseudo-labelings. Figure 3 details the number of samples for each level of unanimity, and Figure 4 relative class frequencies for each level of unanimity. The numerical data displayed in this figure can be found in Supplementary Table IV.
@@
III-C Training with Pseudo-Labeled Data
@@
The cross-validated F1 metrics for utilizing the two sets of pseudo-labeled data according to this strategy are shown in Figure 5. We also chose to include the F1 metrics of RoBERTa trained purely on the 500 labeled posts to serve as a point of comparison. A more detailed report comparing the three frameworks can be found in Supplementary Table V.
@@
III-D Lessons Learned from LLMs
In addition to evaluating the performance of transformer-based models, we also chose to evaluate Llama3-8B as described in Section II-C. Given that Llama3-8B is very time-consuming to fine-tune and that our results indicated Llama3-8B performed worse than RoBERTa, we did not conduct five-fold cross-validation. Therefore the results shown in Figure 6 reflect only one training fold. A more detailed classification report can be found in Supplementary Table VI.
@@ IV Discussion
The results from our experiment, displayed in Section III-C, indicate that self-training with SCS pseudo-label acquisition improves model performance by a substantial margin. Furthermore, the additional step of validating the posts that are not unanimously pseudo-labeled across an ensemble of models provides an additional boost to classification accuracy; however, the majority of the improvement in model performance comes from the addition of the pseudo-labeled data, not the partial-validation. While our proposed SCS strategy exhibits empirical success, it is important to note that our method relies on the assumption that the class distribution in the unlabeled data is equivalent or near-equivalent to that of the labeled data, since the class distribution of the model’s predictions, and thus the distribution of our pseudo-labeled data, is closely tied to the class distribution in our initial dataset. Future deployments of this method should be mindful of this assumption.
@@
Additionally, the data we record in Section III-C suggests that self-training with SCS pseudo-label acquisition disproportionately boosts F1 scores in classes underrepresented in the initial dataset. This indicates that, in addition to addressing the shortcomings presented by traditional pseudo-label selection techniques when dealing with imbalanced datasets, our SCS strategy is also able to partially correct the disparity in model performance between imbalanced classes. Therefore, our SCS strategy also presents a promising method for addressing class-imbalanced partially labeled tasks beyond just suicide risk assessment that we urge future researchers to validate and explore. This effect is even more pronounced for RoBERTa trained on both the labeled and partially validated pseudo-labeled data. Section III-B details the class distribution of the unanimous versus the non-unanimous pseudo-labels and demonstrates that the set of “confusing” samples contains a greater percentage of the “Attempt” class than the unanimously labeled posts. Given the large improvement in the F1 score for the “Attempt” class for RoBERTa trained on the corrected pseudo-labels, we believe correcting the pseudo-labels that tend to confuse the pseudo-labeling algorithm presents a way to improve model performance in predicting underrepresented classes, as those underrepresented classes tend to be the ones confusing the model most frequently.
@@
Furthermore, our results in Section III-D compare the performance of Llama3-8B to RoBERTa. While we ultimately found that RoBERTa achieved superior performance and was computationally cheaper, comparing the performance of LLMs to BERT-based models is an important effort to gauge the ability of these models to perform non-generative tasks. Our finding that RoBERTa outperforms Llama3-8B on text-classification tasks is consistent with prior literature [29] and poses an interesting question about the strengths and weaknesses of LLMs. However, given the specific nature of this task, our findings should not be treated as an absolute indicator of Llama3’s text-classification abilities, and we implore future researchers to investigate the viability of LLMs in other text-classification tasks.
@@ Acknowledgments
Supplementary Table I: Prompt for Llama3
@@
Supplementary Table II: Detailed classification report for BERT, BiomedBERT, and RoBERTa trained on 500 labeled instances.
@@
Supplementary Table III: Agreement between 5 models and majority vote pseudo-labels
@@
Supplementary Table IV: Distribution in unanimity and class between five sets of pseudo-labels
@@
Supplementary Table V: Detailed classification report for RoBERTa trained on labeled data, labeled and pseudo-labeled (PL) data, and labeled and partially validated pseudo-labeled (PVPL) data.
@@
Supplementary Table VI: Detailed classification report for Llama3 8-B zero-shot evaluated on the 500 labeled instances as well as Llama3 8-B fine-tuned on the 500 labeled instances for Causal LM and Sequence Classification.

### v2

@@ Abstract
keywords:
Natural Language Processing Suicide Risk Social Media, Semi-Supervised Learning
1 Introduction
@@
2 Materials and methods
2.1 Task and Data Description
The Suicide Detection on Social Media BigData Cup Challenge involves creating a predictive model to classify Reddit posts into four suicide risk levels: 1) Indicator: The post content has no explicit expression concerning suicide; 2) Ideation: The post content has explicit suicidal expression but there is no plan to commit suicide; 3) Behavior: The post content has explicit suicidal expression and a plan to commit suicide or self-harming behaviors; and 4) Attempt: The post content has explicit expressions concerning historic suicide attempts.
As shown in Table 1, the challenge organizers developed a corpus composed of 500 labeled and 1,500 unlabeled instances. A test dataset that includes 100 unlabeled instances is also included [12]. Each instance contains the text content of a Reddit post taken from the r/SuicideWatch sub-Reddit.
However, two major challenges are presented with this dataset. The first challenge is that only a small portion of available data is labeled, limiting the efficacy of a traditional supervised approach. The second challenge is that the intrinsic class-imbalance present in the labeled dataset makes it difficult for models trained on the provided data to predict classes underrepresented in the labeled corpus (i.e., Attempt).
Table 1: Statistics of the Suicide Detection on Social Media dataset.
@@ 2.2 Method Development
2.2.1 Backbone Model Selection
@@
2.2.2 Semi-Supervised Learning with Stratified Confidence Sampling
@@ 2.2.3 Label Correction
2.3 LLMs with Zero-Shot Learning and Fine-Tuning
@@
We evaluated the capability of Llama3-8B to detect suicide risk levels in the zero-shot learning setting and fine-tuning setting. For fine-tuning, we compared two strategies: next-token prediction and sequence classification. For both zero-shot learning and fine-tuning, the prompt starts with the definitions of the four suicide risk levels, followed by the input text (see Supplementary Table 1).
2.4 Evaluation Metrics
@@ 2.5 Experimental Settings
3 Results
3.1 RoBERTa Outperforms Other Models
Before evaluating our semi-supervised approach, we first fine-tuned three transformer-based models using a supervised approach to determine which embedding model performs the best and establish a baseline against which we can compare our semi-supervised methods. Three transformer-based models, BERT [16], BiomedBERT [17], and RoBERTa [14] were trained and validated according to our experimental settings outlined in Section 2.5. Figure 2 shows the five-fold cross-validated performance of our systems. In columns labeled ”Indicator”, ”Ideation”, ”Behavior”, and ”Attempt”, we plot the binary-F1 score corresponding to posts of that class. The F1 scores yielded by our supervised models indicate that RoBERTa outperformed the other two embedding models across all F1 metrics (both class-specific and multi-class). This improvement is especially pronounced when examining the class-specific metrics. A more detailed report comparing the performance between the three models can be found in Supplementary Table 2.
@@ 3.1 RoBERTa Outperforms Other Models
After deciding on RoBERTa as an embedding model, we conducted self-training with SCS pseudo-label acquisition to generate a dataset of 1,500 pseudo-labeled posts according to the methods outlined in Section 2.2.2. Furthermore, we also partially verified these pseudo-labels according to the strategy outlined in Section 2.2.3.
3.2 Correcting Initial Pseudo-Labels
@@
After training five models and generating five sets of pseudo-labels, we calculated the agreement between each set of pseudo-labels and the majority-vote pseudo-labels using Cohen’s Kappa score [28] (Supplementary Table 3). The rate of agreement between the five sets of pseudo-labels is high, which makes sense given that the training data between any two of the five models is 75% similar. To determine which samples to annotate, the final set of 1,500 pseudo-labeled posts was first split into four separate subsets based on the number of votes the majority voted class received between the 5 initial pseudo-labelings. Figure 3 details the number of samples for each level of unanimity, and Figure 4 relative class frequencies for each level of unanimity. The numerical data displayed in this figure can be found in Supplementary Table 4.
@@
3.3 Training with Pseudo-Labeled Data
@@
The cross-validated F1 metrics for utilizing the two sets of pseudo-labeled data according to this strategy are shown in Figure 5. We also chose to include the F1 metrics of RoBERTa trained purely on the 500 labeled posts to serve as a point of comparison. A more detailed report comparing the three frameworks can be found in Supplementary Table 5.
@@
3.4 Lessons Learned from LLMs
In addition to evaluating the performance of transformer-based models, we also chose to evaluate Llama3-8B as described in Section 2.3. Given that Llama3-8B is very time-consuming to fine-tune and that our results indicated Llama3-8B performed worse than RoBERTa, we did not conduct five-fold cross-validation. Therefore the results shown in Figure 6 reflect only one training fold. A more detailed classification report can be found in Supplementary Table 6.
@@ 4 Discussion
The results from our experiment, displayed in Section 3.3, indicate that self-training with SCS pseudo-label acquisition improves model performance by a substantial margin. Furthermore, the additional step of validating the posts that are not unanimously pseudo-labeled across an ensemble of models provides an additional boost to classification accuracy; however, the majority of the improvement in model performance comes from the addition of the pseudo-labeled data, not the partial-validation. While our proposed SCS strategy exhibits empirical success, it is important to note that our method relies on the assumption that the class distribution in the unlabeled data is equivalent or near-equivalent to that of the labeled data, since the class distribution of the model’s predictions, and thus the distribution of our pseudo-labeled data, is closely tied to the class distribution in our initial dataset. Future deployments of this method should be mindful of this assumption.
@@
Additionally, the data we record in Section 3.3 suggests that self-training with SCS pseudo-label acquisition disproportionately boosts F1 scores in classes underrepresented in the initial dataset. This indicates that, in addition to addressing the shortcomings presented by traditional pseudo-label selection techniques when dealing with imbalanced datasets, our SCS strategy is also able to partially correct the disparity in model performance between imbalanced classes. Therefore, our SCS strategy also presents a promising method for addressing class-imbalanced partially labeled tasks beyond just suicide risk assessment that we urge future researchers to validate and explore. This effect is even more pronounced for RoBERTa trained on both the labeled and partially validated pseudo-labeled data. Section 3.2 details the class distribution of the unanimous versus the non-unanimous pseudo-labels and demonstrates that the set of “confusing” samples contains a greater percentage of the “Attempt” class than the unanimously labeled posts. Given the large improvement in the F1 score for the “Attempt” class for RoBERTa trained on the corrected pseudo-labels, we believe correcting the pseudo-labels that tend to confuse the pseudo-labeling algorithm presents a way to improve model performance in predicting underrepresented classes, as those underrepresented classes tend to be the ones confusing the model most frequently.
@@
Furthermore, our results in Section 3.4 compare the performance of Llama3-8B to RoBERTa. While we ultimately found that RoBERTa achieved superior performance and was computationally cheaper, comparing the performance of LLMs to BERT-based models is an important effort to gauge the ability of these models to perform non-generative tasks. Our finding that RoBERTa outperforms Llama3-8B on text-classification tasks is consistent with prior literature [29] and poses an interesting question about the strengths and weaknesses of LLMs. However, given the specific nature of this task, our findings should not be treated as an absolute indicator of Llama3’s text-classification abilities, and we implore future researchers to investigate the viability of LLMs in other text-classification tasks.
@@ Acknowledgments
Supplementary Table 1: Prompt for Llama3
@@
Supplementary Table 2: Detailed classification report for BERT, BiomedBERT, and RoBERTa trained on 500 labeled instances.
@@
Supplementary Table 3: Agreement between 5 models and majority vote pseudo-labels
@@
Supplementary Table 4: Distribution in unanimity and class between five sets of pseudo-labels
@@
Supplementary Table 5: Detailed classification report for RoBERTa trained on labeled data, labeled and pseudo-labeled (PL) data, and labeled and partially validated pseudo-labeled (PVPL) data.
@@
Supplementary Table 6: Detailed classification report for Llama3 8-B zero-shot evaluated on the 500 labeled instances as well as Llama3 8-B fine-tuned on the 500 labeled instances for Causal LM and Sequence Classification.